Probability Thinking

Part A: Find the number of outcomes in the complement of the given event. Out of 292 high school students in their freshman year, 189 are right-handed.

Part B: User passwords for a certain computer network consist of 3 letters followed by 4 numbers. How many different passwords are possible?

Part C: There are 16 people in an office with 6 different phone lines. If all the lines begin to ring at once, how many groups of 6 people can answer these lines?
Answer #1

Solution

(a) P(A) = 189/292 = 0.6473

P(A') = 1- P(A) = 1-0.6473 = 0.3527

(b) Number of passwords possible = (if letters and numbers cannot be repeated) = 26*25*24*10*9*8*7 = 78624000

(c) 16C6 people can answer these lines i.e. = 16!\(6!10!) = 8008